find the area of the rectangle
One side is a square root of 3 + Square root of 5
and the other side is square root of 3 + Square root of 5

Answer by stanbon(48535)

One side is a square root of 3 + Square root of 5
and the other side is square root of 3 + Square root of 5
----------
(sqrt3 + sqrt5)^2 = 3 + 2sqrt15 + 5 = 8 + 2sqrt15
==============
Cheers,
Stan H.